Let ܺ denote the number of scores on a test. If ܺ is normally

distributed with mean 100 and variance 225, find the probability

that ܺ does not exceed 130.


1
Expert's answer
2021-05-07T11:58:40-0400

"\sigma" = "\sqrt{\sigma^2}" = "\sqrt{225}" =15

The probability that X < 30 is equal to the blue area under the curve.



Since "\mu=100" and "\sigma=15" we have:

P(X<130) = P(X-"\mu" < 130-100) = P"(\frac{X-\mu}{\sigma}<\frac{130-100}{15})"

Since "\frac{X-\mu}{\sigma} = Z" and "\frac{130-100}{15} = 2"  we have:

P(X<130) = P(Z<2)

Use the standard normal table



 to conclude that:

P(Z<2) = 0.9772

Answer: P = 0.9772.